# Disentangling the galactic and intergalactic components in 313 observed Lyman-alpha line profiles between redshift 0 and 5

### v1

@@ Appendix A $f_{\rm esc}^{4\AA}$ inference by an alternative model
We find the same trends in the performance of IGM+z, IGM-z, and REC+z in the observed spectrum as the mock spectrum. The distribution of the predicted $f_{\rm esc}^{4\AA}$ values of the observed ${\rm Ly}\alpha$ line profiles presented in Sect. 4 for REC+z (purple), IGM+z (blue,) and IGM-z (green) are shown in Fig. 16. The left panel shows HST sources at $z<0.55$, while the right panel shows MUSE sources at $z>2.9$. For MUSE sources, the three models show very similar $f_{\rm esc}^{4\AA}$ distributions with $\langle f_{\rm esc}^{4\AA}\rangle$ $\sim 0.7$ (horizontal dashed lines). Not only distribution in $6.02.9<z<$ 6.0 is the same, but also the redshift evolution, although it is not explicitly shown. Meanwhile, considering HST sources, the IGM+z and IGM-z exhibit very similar $f_{\rm esc}^{4\AA}$ distribution, with $\langle f_{\rm esc}^{4\AA}\rangle$ $\sim 0.89$ and $0.91$ respectively. Meanwhile, REC+z predicts higher $f_{\rm esc}^{4\AA}$ values, as the $f_{\rm esc}^{4\AA}$ distribution peaks at $\sim 1$ with small dispersion. Actually, the minimum $f_{\rm esc}^{4\AA}$ predicted by REC+z is $\sim 0.9$. Consequently, the $\langle f_{\rm esc}^{4\AA}\rangle$ for REC+z is $\sim 0.98$.
@@
The training set in an artificial neural network determines its output fidelity. In particular, biases and trends inherent to the training set are imprinted on the output. We find that REC+z gives a $\langle f_{\rm esc}^{4\AA}\rangle$ of $\sim 1$ at $z<0.5$ in contrast to IGM+z and IGM-z that gives $\langle f_{\rm esc}^{4\AA}\rangle$ $\sim 0.9$. Meanwhile, the three models give very similar predictions at high redshift. Therefore, the low $z$ discrepancy comes from the training set of REC+z. The low $f_{\rm esc}^{4\AA}$ variance at $z<1$ in REC+z causes its predictions to be biased towards $f_{\rm esc}^{4\AA}$ =1 at low $z$, rather than being a real prediction.
Appendix B Property correlation in MUSE and HST
In Fig.17 we show the show the spearman correlation coefficients for the HST (left) and MUSE (right) samples. We find that in general, all the correlations found in the complete sample (see Fig.8) are also present in the individual samples, such as the correlation between $EW_{\rm in}$ and $EW_{\rm ob}$.
Other correlations existing in the combined sample are only present in one of them. For example, the anti correlation between $f_{\rm esc}^{4\AA}$ and redshift is present in the MUSE and combined sample. However, no trend is found within the HST sources.
Appendix C all fitted lines
@@ Appendix B Property correlation in MUSE and HST
In this appendix, we show all the fitted lines for both MUSE and HST data. Line profiles are sorted by redshift. Fig. 18 shows sources with $0.200.0<z<$ 0.20, Fig. 19 for sources with $0.350.20<z<$ 0.35, Fig. 20 for sources with $3.30.35<z<$ 3.3 , Fig. 21 for sources with $3.73.3<z<$ 3.7 , Fig. 22 for sources with $4.23.7<z<$ 4.2 ,Fig. 23 for sources with $4.84.2<z<$ 4.8 and Fig. 24 shows sources with $6.04.8<z<$ 6.0. The observed ${\rm Ly}\alpha$ line profile is shown in grey, while the prediction of IGM-z, IGM+z, and NoIGM are shown in blue, green, and yellow, respectively. Furthermore, $f_{\rm esc}^{4\AA}$ is shown in the top left corner in the same color as the model used, alongside the source redshift. The grey shaded band marks the wavelength region where $f_{\rm esc}^{4\AA}$ is estimated.
@@
In Fig. 25, we show the ${\rm Ly}\alpha$ line profiles from LASD that were not considered for the analysis in this work. They removed these line profiles because i) they looked noisy or ii) they exhibited a very steep continuum not modeled in zELDA, or ii) the fits of IGM-z, IGM+z, and NoIGM looked unsuccessfully. From the original sample of 349 spectra, 35 (10%) were excluded.
@@
In general, at low redshift (Fig. 18 and 19) IGM+z, IGM-z and NoIGM predict similar intrinsic ${\rm Ly}\alpha$ line profiles. In sources 7, 40 and 111 IGM+z and IGM-z predict almost the same ${\rm Ly}\alpha$ line profile and NoIGM predicts a different one. In sources 1, 14, 30, 38, 66, 82 and 97 IGM+z and IGM-z actually predict a different intrinsic ${\rm Ly}\alpha$ line profile. Also, in general, the $f_{\rm esc}^{4\AA}$ predicted by IGM+z and IGM-z are close to unity, being IGM-z $f_{\rm esc}^{4\AA}$ predictions slightly lower than those of IGM+z, in general.
@@
Meanwhile, at higher redshift (Fig. 21, 22, 23, and 24), the intrinsic ${\rm Ly}\alpha$ line profile predicted by IGM+z and IGM-z match, while they diverge from that of NoIGM. Although in some cases, the three models match relatively well, such in sources 116, 119, 142, 162, 186, 220, and 299. Note that this scenario is more frequent at lower redshift, closer to $z=3.0$. Meanwhile, at higher redshifts, closer to $z=6.0$, there is more difference between the prediction and IGM+z/IGM-z and NoIGM. Also, in $\sim$80of the high redshift sources, the IGM+z and IGM-z predict the presence of a blue peak in the ${\rm Ly}\alpha$ intrinsic spectrum.
@@
Figure 18: zELDA’s prediction on observed line profiles displayed in restframe. The observed line profile is shown in grey. zELDA’s reconstruction using the models IGM+z, IGM-z, and NoIGM are displayed in blue, green, and yellow, respectively. The redshift of the source is shown in the top left corner. zELDA’s estimations of $f_{\rm esc}^{4\AA}$ given by the IGM+z and IGM-z models are shown in blue and green, respectively. The grey-shaded region shows the wavelength interval of $f_{\rm esc}^{4\AA}$. The red number indicate the index in LASD.
@@
Figure 25: Same as Fig. 18, but for other observed ${\rm Ly}\alpha$ line profiles for which exhibit low $S/N_{p}$ (e.g., 0 , 1 ), a steep continuum not modeled in zELDA (e.g., 0 , 1 ), or with for than two components (e.g., 3).

### v2

@@ Appendix E $f_{\rm esc}^{4\AA }$ inference by an alternative model
We find the same trends in the performance of IGM+z, IGM-z, and REC+z in the observed spectrum as in the mock spectrum. The distribution of the predicted $f_{\rm esc}^{4\AA }$ values of the observed ${\rm Ly}\alpha$ line profiles presented in Sect. 4 for REC+z (purple), IGM+z (blue,) and IGM-z (green) are shown in Fig. 18. The left panel shows HST sources at $z<0.55$, while the right panel shows MUSE sources at $z>2.9$. For MUSE sources, the three models show very similar $f_{\rm esc}^{4\AA }$ distributions with $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.7$ (horizontal dashed lines). Not only is the distribution in $2.9<z<6.0$ the same, but also the redshift evolution, although it is not explicitly shown. Meanwhile, considering HST sources, the IGM+z and IGM-z exhibit very similar $f_{\rm esc}^{4\AA }$ distribution, with $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.89$ and $0.91$ respectively. Meanwhile, REC+z predicts higher $f_{\rm esc}^{4\AA }$ values, as the $f_{\rm esc}^{4\AA }$ distribution peaks at $\sim 1$ with small dispersion. Actually, the minimum $f_{\rm esc}^{4\AA }$ predicted by REC+z is $\sim 0.9$. Consequently, the $\langle f_{\rm esc}^{4\AA }\rangle$ for REC+z is $\sim 0.98$.
@@
The training set in an artificial neural network determines the fidelity of its output. In particular, biases and trends inherent to the training set are imprinted on the output. We find that REC+z gives a $\langle f_{\rm esc}^{4\AA }\rangle$ of $\sim 1$ at $z<0.5$ in contrast to IGM+z and IGM-z that give $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.9$. Meanwhile, the three models give very similar predictions at high redshift. Therefore, the low $z$ discrepancy comes from the training set of REC+z. The low $f_{\rm esc}^{4\AA }$ variance at $z<1$ in REC+z causes its predictions to be biased towards $f_{\rm esc}^{4\AA }$ =1 at low $z$, rather than being a real prediction.
Appendix F All fitted lines
@@
In this appendix, we show all the fitted lines for both MUSE and HST data. Line profiles are sorted by redshift. Fig. 19 shows sources with $0.0<z<0.20$, Fig. 20 for sources with $0.20<z<0.35$, Fig. 21 for sources with $0.35<z<3.3$ , Fig. 22 for sources with $3.3<z<3.7$ , Fig. 23 for sources with $3.7<z<4.2$ ,Fig. 24 for sources with $4.2<z<4.8$ and Fig. 25 shows sources with $4.8<z<6.0$. The observed ${\rm Ly}\alpha$ line profile is shown in grey, while the prediction of IGM-z, IGM+z, and NoIGM are shown in blue, green, and yellow, respectively. Furthermore, $f_{\rm esc}^{4\AA }$ is shown in the top left corner in the same color as the model used, alongside the source redshift. The grey shaded band marks the wavelength region where $f_{\rm esc}^{4\AA }$ is estimated.
@@
In Fig. 26, we show the ${\rm Ly}\alpha$ line profiles from LASD that were not considered for the analysis in this work. They removed these line profiles because i) they looked noisy or ii) they exhibited a very steep continuum not modeled in zELDA, or iii) the fits of IGM-z, IGM+z, and NoIGM looked unsuccessful. From the original sample of 349 spectra, 35 (10%) were excluded.
@@
In general, at low redshift (Fig. 19 and 20) IGM+z, IGM-z and NoIGM predict similar intrinsic ${\rm Ly}\alpha$ line profiles. In sources 7, 40, and 111 IGM+z and IGM-z predict almost the same ${\rm Ly}\alpha$ line profile, and NoIGM predicts a different one. In sources 1, 14, 30, 38, 66, 82, and 97 IGM+z and IGM-z actually predict a different intrinsic ${\rm Ly}\alpha$ line profile. Also, in general, the $f_{\rm esc}^{4\AA }$ predicted by IGM+z and IGM-z are close to unity, with IGM-z $f_{\rm esc}^{4\AA }$ predictions slightly lower than those of IGM+z.
@@
Meanwhile, at higher redshift (Fig. 22, 23, 24, and 25), the intrinsic ${\rm Ly}\alpha$ line profile predicted by IGM+z and IGM-z match, while they diverge from that of NoIGM. Although in some cases, the three models match relatively well, such as in sources 116, 119, 142, 162, 186, 220, and 299. Note that this scenario is more frequent at lower redshift, closer to $z=3.0$. Meanwhile, at higher redshifts, closer to $z=6.0$, there is more difference between the prediction and IGM+z/IGM-z and NoIGM. Also, in $\sim$80% of the high redshift sources, the IGM+z and IGM-z predict the presence of a blue peak in the ${\rm Ly}\alpha$ intrinsic spectrum.
@@
Figure 19: zELDA’s prediction on observed line profiles displayed in restframe. The observed line profile is shown in grey. zELDA’s reconstruction using the models IGM+z, IGM-z, and NoIGM is displayed in blue, green, and yellow, respectively. The redshift of the source is shown in the top left corner. zELDA’s estimations of $f_{\rm esc}^{4\AA }$ given by the IGM+z and IGM-z models are shown in blue and green, respectively. The grey-shaded region shows the wavelength interval of $f_{\rm esc}^{4\AA }$. The red number indicates the index in LASD.
@@
Figure 26: Same as Fig. 19, but for other observed ${\rm Ly}\alpha$ line profiles that exhibit low $S/N_{p}$ (e.g., 0 , 1 ), a steep continuum not modeled in zELDA (e.g., 0 , 1 ), or with more than two components (e.g., 3).